# Shared Spatial Memory Through Predictive Coding

### v1

@@ Results
Beyond single-neuron effects, the specialized units form a population code for higher-order relational variables, most notably inter-agent distance. We identify subpopulations selectively tuned to near, mid-range, and far separations (Fig. 4c). Their graded tuning curves spanned the full range of separations encountered during exploration, yielding a tiling-like coverage of relational distance space and supporting a continuous representation of proximity to the partner (Fig. 4c, bottom right). This organization is consistent with the view that predictive objectives shape compact, task-relevant embeddings of spatial relations [47, 24].
@@
We then quantitatively dissociate these populations using mutual information (MI) between firing rate and spatial variables. Place cells carry high MI about self-position but negligible MI about partner-position, whereas SPCs show the inverse profile (Fig. 4d, top). A scatter of Self MI versus Partner MI reveals clearly separable clusters under our metrics, corresponding to self-tuned, partner-tuned, and mixed-selectivity units (Fig. 4d, bottom). These results indicate that specialization is not confined to a small number of idiosyncratic units but reflects a broader division of labor induced by the social predictive objective.
@@
To assess causal involvement, we perform in-silico lesions. Targeted lesioning of distance-tuned SPCs produced a marked, selective impairment in inter-agent distance prediction relative to pre-lesion performance and to size-matched random lesions (Fig. 4e). A broader lesion of the peer-processing module degrades general position prediction. Meanwhile, these patterns indicate that SPCs are causally important for computing relational social geometry in our setting, whereas self-tuned units primarily support self-localization. Control ablations that preserve overall parameter count but disrupt SPC-selective pathways yielded similar deficits in distance estimation (Fig. 4e).
@@
We next link representation to learning dynamics and task performance. The fully trained network’s ability to predict a partner’s future trajectory depends on the integrity of SPCs: lesioning these units reduces predictive accuracy toward untrained levels (Fig. 4f, left). During training, validation loss decreases as the fraction of specialized units increases, with self-tuned and partner-tuned populations emerging in parallel (Fig. 4f, right). This co-evolution suggests that a functionally segregated social place code is not incidental but emerges as a principal mechanism by which the model solves partner-state prediction and coordination in our framework.
Finally, we clarify scope and relation to prior work. The present results do not claim a novel physiological class. Rather, they show that under a social predictive objective, artificial networks reproduce a division of labor analogous to self- and partner-centered spatial coding, and this division carries demonstrable functional and causal relevance for constructing and exploiting a shared spatial memory. In combination with prior evidence that predictive pressures can organize grid- and distance-related codes in artificial agents [47, 40, 24], these findings support the view that minimizing mutual predictive uncertainty induces specialized representations for self and partner that jointly enable robust coordination under bandwidth and observability constraints.
@@
The preceding sections demonstrate how a unified predictive objective forges specialized components for perception, communication, and social representation. Here, we evaluate the performance of the fully integrated system to validate its collective effectiveness in solving complex, communication-constrained tasks. We assess the complete hierarchical reinforcement learning with intrinsic curiosity module (HRL-ICM) against strong baselines in the Memory-Maze benchmark (Fig. 5a). This framework implements the third level of our predictive coding hierarchy. It uses a policy guided by predictive uncertainty, where an intrinsic reward generated from prediction error directs agents to explore regions that maximally reduce their world model’s uncertainty.
We first assess the framework’s effectiveness across a wide range of navigation problems. In a large-scale test over 10,000 procedurally generated random mazes, HRL-ICM achieves the highest task success rate (72.4%) while simultaneously requiring the fewest steps on average to locate the goal (Fig. 5b). This strong average-case performance is mirrored in environments designed to probe specific aspects of collective intelligence. In a central coordination maze that demands efficient division of labor, our method again achieves the highest success rate of 72.0% (Fig. 5c). Furthermore, in a deceptive maze with numerous long dead-ends, the HRL-ICM agents maintain a high success rate of 66.0% (Fig. 5d). Besides, Supplementary Video 1 directly visualizes the performance comparison in memory maze.This resilience highlights the functional value of the emergent communication mechanism, which enables agents to share high-value negative information, a critical capability for efficient exploration.
@@
A critical test for any multi-agent system is scalability. We evaluate how the framework performs as the number of agents increased from two to five (Fig. 5f). Our HRL-ICM framework demonstrates high scalability, maintaining a stable success rate across varying numbers of agents. In contrast, baseline strategies that rely on naive broadcasting (Full and Periodic) suffers a clear performance decline as more agents are added. Their inability to manage the exponential growth in information flow leads to channel congestion and degraded coordination. Our framework’s learned, predictive communication mechanism mitigates this issue by ensuring that only the most vital, non-redundant information is transmitted.
Finally, we test the framework’s robustness under its core design constraint: limited communication bandwidth (Fig. 5g). The performance heatmap reveals the system’s high resilience to bandwidth restrictions. As the available bandwidth is reduced from 128 bits/step to an extremely constrained 4 bits/step, HRL-ICM’s success rate degrades only moderately from 73.5% to 64.4%. This contrasts sharply with the Full Broadcast baseline, whose performance collapsed under the same conditions, plummeting from 67.6% to 28.6%. This result provides compelling evidence for the efficacy of our approach. By learning to communicate only the essential prediction errors of their partners, the agents can sustain a high degree of coordinated action in austere communication environments where conventional methods fail. This demonstrates a robust solution to the challenge of communication-limited coordination.
@@
Having demonstrated the functional benefits of the integrated system, we next analyze its learning dynamics, the causal nature of its emergent communication, and its ability to generalize. A powerful internal model is only useful if it arises from a stable learning process, translates into effective behavior, and scales to challenges beyond its training distribution.
@@
A prerequisite for any complex autonomous system is convergence to a stable, high-performance policy. We demonstrate this property by tracking the training dynamics of our framework across mazes of varying difficulty (Fig. 6a). In all tested environments, from $15\times 15$ to $39\times 39$ configurations, both the task success rate and the average reward per episode exhibit a steady and robust increase, ultimately converging to stable plateaus. This validates our end-to-end training approach. It also shows that the global objective of minimizing predictive uncertainty provides a consistent and effective learning signal for both visuomotor control and inter-agent communication. The learning process is not brittle but reliably guides agents toward a competent collaborative strategy.
While the emergent communication mechanism is efficient, a critical question remains: Do the learned symbols possess a grounded, causal meaning shared between agents? To disambiguate this, we designed a causal intervention experiment (Fig. 6b). In a control scenario, a sender agent correctly identifies a target and transmits a corresponding message to a receiver, which then successfully navigates to the correct location. We then perform a counterfactual intervention by intercepting the true message and replacing it with a “fake message”, a symbol previously associates with a distractor’s location. The effect is immediate and unambiguous. Upon receiving the fake message, the receiver predictably changes its trajectory and navigates directly to the fake target. This specific change in behavior in response to a manipulated signal provides powerful evidence that the learned symbols are not mere correlations but function as causal drivers of the receiver’s actions. This confirms that the agents develop a shared, grounded understanding of the emergent symbols, forming the basis for language-driven coordination. Moreover, the causal intervention experiment is visualized in Supplementary Video 3, which demonstrates the receiver agent’s trajectory being predictably altered by the manipulated message.
@@
The ultimate test of the framework lies in its ability to generalize its learned strategies to environments more complex than those seen during training. We evaluate this by comparing the HRL-ICM framework against baselines across four scales of maze complexity, from $25\times 25$ to $39\times 39$ (Fig. 6c). The results demonstrate both superior performance and scalability. Across all maze sizes, our framework achieves a significantly higher task success rate (Fig. 6d) and requires fewer steps for completion (Fig. 6e) than all baselines. Critically, as environmental complexity increased, the performance of the baseline strategies degrades sharply. In contrast, our framework exhibites a much more graceful decline in performance. This superior scalability is a direct consequence of the framework’s core principles. Instead of relying on brute-force information sharing, agents in our framework leverage their learned internal models, such as the grid-cell-like metric scaffold and social place cell code, to form a robust shared memory. They use the emergent communication mechanism to transmit only the most critical, uncertainty-reducing information. This is an inherently efficient strategy that is less susceptible to the combinatorial explosion of the state space. These findings validate that our theoretically-grounded approach resolves the trade-off between communication bandwidth and coordination effectiveness, yielding a generalizable and scalable solution for multi-agent intelligence.
@@ Discussion
We propose that a single computational objective—the minimization of mutual predictive uncertainty—can serve as a first principle for the emergence of collective intelligence. We demonstrate that this unified predictive drive gives rise to a three-tiered hierarchy of phenomena: a stable, grid-cell-like spatial metric at the individual level; an efficient communication protocol at the inter-actional level; and a specialized neural substrate for social cognition, analogous to hippocampal social place cells, at the representational level. Our findings suggest that complex neural architectures supporting social navigation may not be distinct adaptations but are the computational consequence of this predictive learning framework. This frames social intelligence as an extension of individual cognition.
@@
Our work is grounded in established principles of individual cognition. We first confirm that a stable personal spatial representation, scaffolded by an emergent, grid-cell-like metric critical for robust navigation [2], is a necessary foundational component (Fig. 2). While architectures such as the Tolman–Eichenbaum Machine provide a unifying account of how an individual agent learns and generalizes spatial and relational knowledge [37], our work addresses the subsequent challenge: how multiple agents, each possessing such internal models, achieve collective intelligence. We propose that the drive to minimize mutual uncertainty provides a computational bridge from individual cognition to shared understanding.
@@
From this foundation, we address emergent communication. Unlike frameworks exploring asymmetric, teacher–student knowledge transfer [39], our setting resolves symmetric, peer-to-peer coordination. Framing communication through the information bottleneck principle, we show that agents can autonomously learn a protocol that balances communication cost and predictive utility (Fig. 3c), surpassing standard multi-agent reinforcement learning baselines, such as PPO and Dreamer, considered here [25]. A key consequence of this social predictive objective is the spontaneous formation of a functionally specialized neural substrate for social cognition. Inspired by the discovery of hippocampal social place cells in bats [26], we provide a computational account consistent with their emergence under a social predictive objective. Whereas biological studies observe these neurons, our framework posits that the imperative to predict a partner’s future state encourages the development of functionally specialized units tuned to the location of others. Our in-silico lesion experiments (Fig. 4e, f) indicate that these emergent “social place cells” are causally important for social prediction, offering a functional interpretation for the “shared neural subspaces” reported in interacting systems [44].
@@
The same principle mirrors the “next-token prediction” objective that underpins large language models (LLMs): while LLMs build a world model by predicting sequences of text, our agents build a shared spatial model by predicting sequences of each other’s sensory states. This parallel suggests that predictive learning is a general mechanism for constructing both individual and shared world models. The connection between understanding and effective compression, highlighted in recent work on lossless data compression [23], provides a theoretical link to the communication efficiency observed in our system. For multi-robot systems, this work offers a paradigm for designing communication-efficient swarms that learn to coordinate from first principles [25].
@@
While our framework establishes these principles, it opens several avenues. The current model utilizes a discrete communication channel; future work could explore the emergence of more continuous or compositional communication structures [39]. Deploying our model-based learning approach on physical multi-robot systems to bridge the simulation-to-reality gap is a key next step [25]. More broadly, the phenomena of shared memory and social representation may arise not from intricate, pre-programmed rules, but from a single computational objective: the drive to predict the world of another.
@@ Methods
Each agent in the multi-agent system operates based on egocentric sensory inputs and is subject to realistic physical and communicational constraints. An agent’s perception is primarily driven by a forward-facing RGB camera providing a 75-degree field of view (FOV), which serves as its visual input. In addition to vision, each agent has access to its own proprioceptive information, namely its linear and angular velocities. The action space is discrete, consisting of move-forward, turn-left, turn-right, stay, enabling navigation through the maze corridors. A critical component of our experimental design is the constraint on inter-agent communication. Agents can only exchange information when they are within a pre-defined communication range, and the channel is subject to a configurable bandwidth bottleneck (e.g., 4 to 128 bits/step). This limitation mirrors real-world robotic applications and creates a strong selective pressure for the emergence of an efficient and targeted communication protocol, which is a central focus of our work. The collective objective for the team of $N$ agents (where $N$ is typically 2-4 in our experiments) is to collaboratively explore the unknown maze to locate a single, hidden target.
@@
An agent must first form a coherent internal model of its own state and surroundings to serve as a foundation for any higher-level reasoning or social interaction. This is achieved not through a single monolithic network, but by the synergistic interplay of two specialized yet deeply integrated sub-systems that solve simultaneous, coupled prediction tasks. Meanwhile, these components function analogously to a visual simultaneous localization and mapping (SLAM) system, where the agent concurrently builds a map of its environment (perception) while predicting its location within it (localization). This dual predictive process is essential because each task regularizes the other: a stable location estimate prevents the map from becoming distorted, while a coherent map provides the landmarks necessary for accurate localization. This reciprocal relationship enables the construction of a reliable individual world model, which is the prerequisite for forming a shared spatial memory.
@@
As shown in Fig. 1b, the primary perceptual task is framed as a visual prediction problem that addresses the challenge of inferring a stable world representation from a fleeting, ambiguous sensory stream. Specifically, the agent must learn a generative model of the physical world’s local geometry and appearance. It does so by learning to translate a high-dimensional, egocentric image, $O_{\text{Ego},t}\in\mathbb{R}^{3\times H_{\text{in}}\times W_{\text{in}}}$ at time $t$, into a structured, allocentric BEV map, $\hat{O}_{\text{BEV},t}\in\mathbb{R}^{4\times H_{\text{out}}\times W_{\text{out}}}$. This is a fundamentally ill-posed inverse problem, and the network is trained end-to-end to minimize the prediction error between its generated map and the ground truth. This error is quantified by a composite loss function, $\mathcal{L}_{\text{BEV}}$, which holistically evaluates the quality of the prediction across multiple modalities. Each component of this loss can be interpreted as imposing a different physical prior on the generative model, guiding it towards physically plausible solutions:
@@
where $w_{\text{occ}}$, $w_{\text{rgb}}$, and $w_{\text{smooth}}$ are weighting hyper-parameters. The occupancy loss, $\mathcal{L}_{\text{occ}}$, is a binary cross-entropy term on the predicted alpha channel ($\hat{\alpha}$) that forces the network to accurately predict the binary state of the world (i.e., navigable or occupied). The appearance loss, $\mathcal{L}_{\text{rgb}}$, a masked mean-squared error term, compels the model to predict the correct target color within navigable regions. Finally, the smoothness loss, $\mathcal{L}_{\text{smooth}}$, regularizes the predictive model by penalizing sharp spatial gradients, incorporating a prior that physical environments are generally continuous. To solve this complex, cross-view prediction task, we employ a sophisticated Transformer-based encoder-decoder architecture. The detailed layer-by-layer specification of this architecture is provided in Supplementary Tables S4 and S5, which leverages a Transformer’s self-attention mechanism to overcome perspective distortion inherent in ground-level views. The encoder, $E_{\text{BEV}}$, uses a ResNet-18 backbone to extract spatial features. A key architectural innovation lies in treating the image’s vertical scanlines as a sequence. This allows the Transformer’s self-attention mechanism to disambiguate visual features by considering the global context along the vertical axis of the image, which is crucial for the predictive model to accurately infer depth and overcome perspective distortion. The decoder, $D_{\text{BEV}}$, then projects the resulting latent code into the final BEV map. The mathematical formulation and weighting of each loss component are detailed in Supplementary Method S1.2, which define the precise objective for learning photometrically and geometrically plausible BEV maps.
@@
The design of this loss function is critical for maintaining a stable predictive process over long trajectories. The primary term, $\mathcal{L}_{\text{KL}}$, applies a significantly higher weight to the prediction error in the initial trajectory frames, forcing the model to establish an accurate pose estimate early to prevent error accumulation. This is supplemented by an initial consistency term, $\mathcal{L}_{\text{init}}$, and a continuity regularizer, $\mathcal{L}_{\text{cont}}$. This structured learning pressure compels the LSTM’s hidden units to discover the most efficient neural code for representing Euclidean space under translational motion. Remarkably, the optimal representation that emerges spontaneously to best solve this temporal self-prediction task is a grid-cell-like code, exhibiting the hexagonal symmetry found in the mammalian entorhinal cortex. This provides strong evidence that such complex, biologically-plausible neural structures can arise not from explicit design, but as a fundamental and convergent solution to the problem of continuous self-prediction in a spatial environment. This emergent metric scaffold provides the essential, stable foundation upon which the local visual predictions can be reliably integrated, forming a coherent and robust individual spatial memory. Theoretically, this learning process allows the LSTM to function as an amortized Bayesian filter, effectively performing probabilistic state estimation, as detailed in Supplementary Method S1.1. The architecture and training details are listed in Supplementary Tables S1 and S2.
@@
Probabilistic formulation. We formalize path integration as a problem of sequential state inference. The agent’s pose at time $t$ is characterized by the state vector $s_{t}=(\mathbf{r}_{t},\theta_{t})\in\mathbb{R}^{2}\times\mathbb{S}^{1}$, where $\mathbf{r}_{t}=(r_{x,t},r_{y,t})^{\top}$ denotes the position and $\theta_{t}$ denotes the heading angle. The agent’s kinematics follow the dynamics:
@@
The path integrator network receives no direct supervision on the latent pose $s_{t}$ during trajectory execution. Instead, its sole learning signal derives from predicting the activity of virtual place and head-direction cells (artificial network units that encode spatial position and heading orientation, respectively, using biologically-inspired activation patterns; see Supplementary Method S1.1 for detailed definitions) at the final timestep. These spatial cells exhibit tuning curves with log-potentials $\phi_{\ell}(s)$: place cells use Gaussian receptive fields $\phi_{i}(\mathbf{r})=-\|\mathbf{r}-\bm{\mu}_{i}\|^{2}/(2\sigma_{i}^{2})$ while head-direction cells employ von Mises tuning $\phi_{j}(\theta)=\kappa_{j}\cos(\theta-\mu_{j})$. The probability of observing cell $\ell$ active in pose $s_{t}$ follows a softmax distribution over these potentials:

### v2

@@ Results
Beyond single-neuron effects, the specialized units form a population code for higher-order relational variables, most notably inter-agent distance. We identify subpopulations selectively tuned to near, mid-range, and far separations (Fig. 4c). Their graded tuning curves span the full range of separations encountered during exploration, yielding a tiling-like coverage of relational distance space (Fig. 4c, bottom right). This organization is consistent with the view that predictive objectives shape compact, task-relevant embeddings of spatial relations [23, 25].
@@
To move beyond descriptive characterization and establish functional relevance, we quantitatively dissociate these populations using mutual information (MI) between firing rate and spatial variables. Place cells carry high MI about self-position but negligible MI about partner-position, whereas SPCs show the inverse profile (Fig. 4d, top). A scatter of Self MI versus Partner MI reveals clearly separable clusters corresponding to self-tuned, partner-tuned, and mixed-selectivity units (Fig. 4d, bottom). These results indicate that specialization reflects a broader division of labor induced by the social predictive objective.
@@
Moreover, to assess causal involvement, we perform in-silico lesions. Targeted lesioning of distance-tuned SPCs produced a marked, selective impairment in inter-agent distance prediction relative to pre-lesion performance and to size-matched random lesions (Fig. 4e). A broader lesion of the peer-processing module degrades general position prediction. These patterns indicate that SPCs are causally important for computing relational social geometry, whereas self-tuned units primarily support self-localization. Control ablations that preserve overall parameter count but disrupt SPC-selective pathways yielded similar deficits in distance estimation.
@@
Finally, linking representation to learning dynamics, the fully trained network’s ability to predict a partner’s future trajectory depends on the integrity of SPCs: lesioning these units reduces predictive accuracy toward untrained levels (Fig. 4f, left). During training, validation loss decreases as the fraction of specialized units increases, with self-tuned and partner-tuned populations emerging in parallel (Fig. 4f, right). This co-evolution suggests that a functionally segregated social place code is not incidental but emerges as a principal mechanism by which the model solves partner-state prediction and coordination. Combined with prior evidence that predictive pressures can organize grid- and distance-related codes in artificial agents [23, 24, 25], these findings support that minimizing mutual predictive uncertainty induces specialized representations for self and partners that jointly enable robust coordination under bandwidth and observability constraints.
@@
The preceding sections demonstrate how a unified predictive objective forges specialized components for perception, communication, and social representation. Here, we assess whether these components synergistically combine to achieve superior cooperative performance. Unlike prior work that addresses individual components in isolation [45, 26], we evaluate the fully integrated system across diverse navigation challenges. We assess the hierarchical reinforcement learning with intrinsic curiosity module (HRL-ICM) against strong baselines in Memory-Maze (Fig. 5a), implementing our three-level predictive hierarchy with policies guided by prediction-error-based intrinsic rewards.
We first assess our framework’s effectiveness across a wide range of navigation problems. Over 10,000 procedurally generated mazes, HRL-ICM achieves highest success rate (72.4%) with fewest steps (Fig. 5b). Moreover, performance holds in targeted scenarios: 72.0% in central coordination mazes (Fig. 5c) and 66.0% in deceptive mazes with long dead-ends (Fig. 5d), highlighting emergent communication’s value for sharing negative information. Additionally, Supplementary Video 1 directly visualizes the performance comparison in the memory maze.This resilience highlights the functional value of the emergent communication mechanism, which enables agents to share high-value negative information, a critical capability for efficient exploration.
@@
A critical test for any multi-agent system is scalability. We evaluate how the framework performs as the number of agents increases from two to five (Fig. 5f). Baseline strategies that rely on simple broadcasting (Full and Periodic) suffer a clear performance decline as more agents are added. Their inability to handle the exponential growth in information flow leads to channel congestion and degraded coordination. The learned, predictive communication mechanism from our framework mitigates this issue by ensuring that only the most vital, non-redundant information is transmitted. Besides, we test bandwidth robustness (Fig. 5g). Reducing bandwidth from 128 to 4 bits/step, HRL-ICM success degrades moderately (12% relative decline) while Full Broadcast collapses (58% decline). By communicating only essential prediction errors, agents sustain coordinated action in austere communication environments where conventional methods fail.
@@
The results presented thus far establish the emergent properties and cooperative performance of our framework. To complete characterization, we address three critical questions: Does learning converge reliably? Do emergent symbols possess genuine semantic meaning? Can learned strategies generalize beyond training conditions?
@@
First, we demonstrate convergence by tracking training dynamics across varying difficulty (Fig. 6a). From $15\times 15$ to $39\times 39$ configurations, success rates and rewards exhibit steady increases, converging to stable plateaus. This validates end-to-end training and confirms minimizing predictive uncertainty provides consistent learning signals for visuomotor control and inter-agent communication.
Second, to test semantic grounding, we designed causal intervention experiments (Fig. 6b). Senders who accurately identify recipients transmit messages that empower receivers. Replacing true messages with “fake” symbols previously associated with distractors causes receivers to predictably alter trajectories toward fake targets. This behavioral change demonstrates learned symbols function as causal drivers, not mere correlations, confirming agents develop shared, grounded understanding. Moreover, the causal intervention experiment is visualized in Supplementary Video 3.
@@
Third, for generalization, we compare HRL-ICM against baselines across four complexity scales (Fig. 6c-e). Our framework achieves significantly higher success rates and requires fewer steps than baselines. Critically, as complexity increases, baseline performance degrades sharply while our framework exhibits graceful decline. This superior scalability stems from core principles: rather than brute-force information sharing, agents leverage learned internal models—grid-cell metric scaffolds and social place codes—forming robust shared memories. The emergent communication mechanism shares only critical, uncertainty-reducing information, an efficient strategy less susceptible to state-space combinatorial explosion. These findings validate our theoretically-grounded approach resolves bandwidth-coordination trade-offs, yielding generalizable, scalable multi-agent intelligence solutions.
@@ Discussion
In this article, we have shown that a single computational objective—the minimization of mutual predictive uncertainty—can serve as a first principle for the emergence of collective intelligence. We demonstrated that this unified predictive drive gives rise to a three-tiered hierarchy of phenomena: a stable, grid-cell-like spatial metric at the individual level; an efficient communication mechanism at the inter-actional level; and a specialized neural substrate for social cognition, analogous to hippocampal social place cells, at the representational level. Our findings suggest that complex neural architectures supporting social navigation may not be distinct adaptations but are the computational consequence of this predictive learning framework, which formulates social intelligence as an extension of individual cognition.
@@
Our work is grounded in established principles of individual cognition. We first confirm that a stable personal spatial representation, scaffolded by an emergent, grid-cell-like metric critical for robust navigation [45], is a necessary foundational component (Fig. 2). While architectures such as the Tolman–Eichenbaum Machine provide a unifying account of how an individual agent learns and generalizes spatial and relational knowledge [19], our work addresses the subsequent challenge: how multiple agents, each possessing such internal models, achieve collective intelligence. We posit that the drive to minimize mutual uncertainty provides a computational bridge from individual cognition to shared understanding.
@@
From this foundation, we address emergent communication. Unlike frameworks exploring asymmetric, teacher–student knowledge transfer [46], our setting resolves symmetric, peer-to-peer coordination. By framing communication through the information bottleneck principle, we show that agents can autonomously learn a communication mechanism that balances communication cost and predictive utility (Fig. 3c), surpassing standard multi-agent reinforcement learning baselines, such as PPO and Dreamer, considered here [47]. A key consequence of this social predictive objective is the spontaneous formation of a functionally specialized neural substrate for social cognition. Inspired by the discovery of hippocampal social place cells in bats [13], we provide a computational account consistent with their emergence under a social predictive objective. Whereas biological studies observe these neurons, our framework posits that the imperative to predict a partner’s future state encourages the development of functionally specialized units tuned to the location of others. Our in-silico lesion experiments (Fig. 4e, f) indicate that these emergent “social place cells” are causally important for social prediction, offering a functional interpretation for the “shared neural subspaces” reported in interacting systems [48].
@@
The same principle mirrors the “next-token prediction” objective that underpins large language models (LLMs): while LLMs build a world model by predicting sequences of texts, our agents build a shared spatial model by predicting sequences of each other’s sensory states. This parallel suggests that predictive learning is a general mechanism for constructing both individual and shared world models. The connection between understanding and effective compression, highlighted in recent work on lossless data compression [49], provides a theoretical link to the communication efficiency observed in our system. For multi-robot systems, this work offers a paradigm for designing communication-efficient swarms that learn to coordinate from first principles [47].
@@
While our framework establishes these principles, it opens several interesting avenues. The current model utilizes a discrete communication channel; future work could explore the emergence of more continuous or compositional communication structures [46]. Deploying our model-based learning approach on physical multi-robot systems to bridge the simulation-to-reality gap is a key next step [47]. More broadly, the phenomena of shared memory and social representation may arise not from intricate, pre-programmed rules, but from a single computational objective: the drive to predict the world of another.
@@ Methods
Each agent in the multi-agent system operates based on egocentric sensory inputs and is subject to realistic physical and communicational constraints. An agent’s perception is primarily driven by a forward-facing RGB camera providing a 75-degree field of view (FOV), which serves as its visual input. In addition to vision, each agent has access to its own proprioceptive information, namely its linear and angular velocities. The action space is discrete, consisting of move-forward, turn-left, turn-right, and stay, enabling navigation through the maze corridors. A critical component of our experimental design is the constraint on inter-agent communication. Agents can only exchange information when they are within a pre-defined communication range, and the channel is subject to a configurable bandwidth bottleneck (e.g., 4 to 128 bits/step). This limitation mirrors real-world robotic applications and creates a strong selective pressure for the emergence of an efficient and targeted communication mechanism, which is a central focus of our work. The collective objective for the team of $N$ agents (where $N$ is typically 2-4 in our experiments) is to collaboratively explore the unknown maze to locate a single, hidden target.
@@
An agent must first form a coherent internal model of its own state and surroundings to serve as a foundation for any higher-level reasoning or social interaction. This is achieved not through a single monolithic network, but by the synergistic interplay of two specialized yet deeply integrated sub-systems that solve simultaneous, coupled prediction tasks. Meanwhile, these components function analogously to a visual simultaneous localization and mapping (SLAM) system, where the agent concurrently builds a map of its environment (perception) while predicting its location within it (localization). This dual predictive process is essential because each task validates the other: a stable location estimate prevents the map from becoming distorted, while a coherent map provides the landmarks necessary for accurate localization. This reciprocal relationship enables the construction of a reliable individual world model, which is the prerequisite for forming a shared spatial memory.
@@
As shown in Fig. 1b, the primary perceptual task is framed as a visual prediction problem that addresses the challenge of inferring a stable world representation from a fleeting, ambiguous sensory stream. Specifically, the agent must learn a generative model of the physical world’s local geometry and appearance. The agent does so by learning to translate a high-dimensional, egocentric image, $O_{\text{Ego},t}\in\mathbb{R}^{3\times H_{\text{in}}\times W_{\text{in}}}$ at time $t$, into a structured, allocentric BEV map, $\hat{O}_{\text{BEV},t}\in\mathbb{R}^{4\times H_{\text{out}}\times W_{\text{out}}}$. This is a fundamentally ill-posed inverse problem, and the network is trained end-to-end to minimize the prediction error between its generated map and the ground truth. This error is quantified by a composite loss function, $\mathcal{L}_{\text{BEV}}$, which holistically evaluates the quality of the prediction across multiple modalities. Each component of this loss can be interpreted as imposing a different physical prior on the generative model, guiding it towards physically plausible solutions:
@@
where $w_{\text{occ}}$, $w_{\text{rgb}}$, and $w_{\text{smooth}}$ are weighting hyper-parameters. The occupancy loss, $\mathcal{L}_{\text{occ}}$, is a binary cross-entropy term on the predicted alpha channel ($\hat{\alpha}$) that forces the network to accurately predict the binary state of the world (i.e., navigable or occupied). The appearance loss, $\mathcal{L}_{\text{rgb}}$, which is a masked mean-squared error term, compels the model to predict the correct target color within navigable regions. Finally, the smoothness loss, $\mathcal{L}_{\text{smooth}}$, regularizes the predictive model by penalizing sharp spatial gradients, incorporating a prior that physical environments are generally continuous. To solve this complex, cross-view prediction task, we employ a sophisticated Transformer-based encoder-decoder architecture. The detailed layer-by-layer specification of this architecture is provided in Supplementary Tables S4 and S5, which leverages a Transformer’s self-attention mechanism to overcome perspective distortion inherent in ground-level views. The encoder, $E_{\text{BEV}}$, uses a ResNet-18 backbone to extract spatial features. A key architectural innovation lies in treating the image’s vertical scanlines as a sequence. This allows the Transformer’s self-attention mechanism to disambiguate visual features by considering the global context along the vertical axis of the image, which is crucial for the predictive model to accurately infer depth and overcome perspective distortion. The decoder, $D_{\text{BEV}}$, then projects the resulting latent code into the final BEV map. The mathematical formulation and weighting of each loss component are detailed in Supplementary Method S2.2, which define the precise objective for learning photometrically and geometrically plausible BEV maps.
@@
The design of this loss function is critical for maintaining a stable predictive process over long trajectories. The primary term, $\mathcal{L}_{\text{KL}}$, applies a significantly higher weight to the prediction error in the initial trajectory frames, forcing the model to establish an accurate pose estimate early to prevent error accumulation. This is supplemented by an initial consistency term, $\mathcal{L}_{\text{init}}$, and a continuity regularizer, $\mathcal{L}_{\text{cont}}$. This structured learning pressure compels the LSTM’s hidden units to discover the most efficient neural code for representing Euclidean space under translational motion. Remarkably, the optimal representation that emerges spontaneously to best solve this temporal self-prediction task is a grid-cell-like code, exhibiting the hexagonal symmetry found in the mammalian entorhinal cortex. This provides strong evidence that such complex, biologically plausible neural structures can arise not from explicit design, but as a fundamental and convergent solution to the problem of continuous self-prediction in a spatial environment. This emergent metric scaffold provides the essential, stable foundation upon which the local visual predictions can be reliably integrated, forming a coherent and robust individual spatial memory. Theoretically, this learning process allows the LSTM to function as an amortized Bayesian filter, effectively performing probabilistic state estimation, as detailed in Supplementary Method S2.1. The architecture and training details are provided in Supplementary Tables S1 and S2.
@@
Probabilistic formulation. We formalize path integration as a problem of sequential state inference. The agent’s pose at time $t$ corresponds to the state vector $s_{t}=(\mathbf{r}_{t},\theta_{t})\in\mathbb{R}^{2}\times\mathbb{S}^{1}$, where $\mathbf{r}_{t}=(r_{x,t},r_{y,t})^{\top}$ denotes the position and $\theta_{t}$ denotes the heading angle. The agent’s kinematics follow the dynamics:
@@
The path integrator network receives no direct supervision on the latent pose $s_{t}$ during trajectory execution. Instead, its sole learning signal derives from predicting the activity of virtual place and head-direction cells (artificial network units that encode spatial position and heading orientation, respectively, using biologically-inspired activation patterns; see Supplementary Method S2.1 for detailed definitions) at the final timestep. These spatial cells exhibit tuning curves with log-potentials $\phi_{\ell}(s)$: place cells use Gaussian receptive fields $\phi_{i}(\mathbf{r})=-\|\mathbf{r}-\bm{\mu}_{i}\|^{2}/(2\sigma_{i}^{2})$ while head-direction cells employ von Mises tuning $\phi_{j}(\theta)=\kappa_{j}\cos(\theta-\mu_{j})$. The probability of observing cell $\ell$ active in pose $s_{t}$ follows a softmax distribution over these potentials: